OpenStudy (anonymous):

Multiply and then simplify the radical 5square root of 10 times 3 square root of 8

OpenStudy (anonymous):

15 square root of 80 then simplify the 80 under the radical \[\sqrt{16*5}\] \[15*4\sqrt{5}\]

OpenStudy (anonymous):

wheres the sqaure root sign on your computer?

OpenStudy (anonymous):

so is it www.dragon.com?

OpenStudy (anonymous):

cAN you give me the url to the website?

OpenStudy (radar):

\[5\sqrt{10}\times 3\sqrt{8} =15\sqrt{80}=15 \times4\sqrt{5}=60\sqrt{5}\]

OpenStudy (anonymous):

can someone please give me the url to dragon.com? or tell me how they get the sqaure root signs on the computer?

OpenStudy (anonymous):

use the Equation button below